Please help me solve the linear system
9x - 7y = 31
-9x + 3y =-39

Answer by macston(5194)

9x - 7y = 31
-9x + 3y =-39 Add the equations.
-4y=-8 Divide each side by -4.
y=2 ANSWER 1: y=2
9x-7y=31 Plug in y value and solve for x
9x-7(2)=31
9x-14=31 Add 14 to each side.
9x=45 Divide each side by 9
x=5 ANSWER 2: x=5
CHECK:
-9x+3y=-39
-9(5)+3(2)=-39
-45+6=-39
-39=-39